A bucket of capacity 10 litres weighs the same as 2 litres of water when
empty. The bucket is inverted and held below the surface of a deep
pond, trapping all the air. Below what depth will the bucket sink of its
own accord? Assume that the bucket, air and pond are all at the same
temperature; that the inverted position is stable; and that the material of
the bucket has negligible volume. You may also assume that one
atmosphere of pressure is equivalent to a 10 metre depth of water.
1
Expert's answer
2020-05-08T16:15:07-0400

The mass of the empty bucket is 2 kg. According to Newton's second law, bucket floats when the force of gravity equals the buoyancy force:


"mg=\rho gV_2,"


where "m" - mass of the bucket, "\rho" - density of water.

This corresponds to the volume of the air bubble of


"V_2=\frac{mg}{\rho g}=\frac{m}{\rho},"


When the volume of the air bubble becomes smaller than this value, the bucket sinks.

The initial pressure of the air bubble in the bucket is the atmospheric pressure "p_1". As the bucket goes deeper, the volume decreases because the water pressure compresses it. Calculate what pressure can compress the air bubble from "V_1=10 \text{ L}" to V assuming the temperature does not change:


"p_1V_1=p_2V_2,\\
p_2=p_1\frac{V_1}{V_2}=p_1\frac{V_1\rho}{m}."

At what depth will the air bubble experience pressure "p_2"? The answer is


"p_2=\rho gh.\\
\rho gh=p_1\bigg(\frac{V_1\rho}{m}\bigg),\\
\space\\
h=\frac{p_1}{\rho g}\bigg(\frac{V_1\rho}{m}\bigg)=\\
\space\\
=\frac{101325}{1000\cdot9.8}\bigg(\frac{10^{-3}\cdot1000}{2}\bigg)=5.2\text{ m}."